4. The time, t, that a butterfly lives after emerging from its chrysalis can be modeled by the probability function 36 20 0, otherwise f(t) , where t is measured in days. a. Sketch the graph of f(t). (What are the x- and y axes labels?) b. Shade the region on your graph that represents the probability that a butterfly will die within 7 days of emerging. c. Compute the probability that a butterfly will die within 7 days of emerging. d. If a large number of butterflies emerge on the same day, after how many days would you expect only 7% to be alive?
Answer #1

Note that expression of f(t) is not correct since 00 36 0 o (6+t)2 So I have modified it as follows 6 (6+ t)2 0 otherwise

a. R code:

t=seq(0,10,by=0.01)
f=6/(6+t)^2
plot(t,f,lwd=2,col=1,type="l",ylab="f(t)",xlab="t")

8 6 4 2 0

b.

(c)

= 6*(1/6-1/(6+7)) = 0.5384

CDF of T is t 36 Fr(t) = P(T < t) = | f(t)dt 6(6+t 6)t o (6+t)2 = 6(1/6-1/(6 + t)) =-6(6 + t) G+ t Now, 10.07 or, t (6t)0.07 or, 0.93t 1 0.07) or t = 6 * (1-0.07)/093 6 Answer 6 days.